OpenStudy (peterg2001):

help please will medal If the diameter of the circle is 11 inches, then the radius will be _____ inches. 22 11 5.5 @k_lynn

OpenStudy (peterg2001):

@Nurali

OpenStudy (k_lynn):

The radius is always half of the diameter. So divide 11 by 2.

OpenStudy (k_lynn):

\[11\div2\]

OpenStudy (anonymous):

the circumference of a circle is equal to 2 * pi * r pi is a constant those value is 3.1415926535898. if the radius is 11 inches, then the circumference is equal to 2 * pi * 11 = 2 * 3.1415926535898 * 11 = 69.1150383789756. you can also leave the answer in terms of pi. in that case, the circumference is equal to 2 * pi * 11 = 22 * pi.

OpenStudy (k_lynn):

Once you divide those two numbers, you get the radius.

OpenStudy (peterg2001):

last one?

OpenStudy (anonymous):

yep

OpenStudy (k_lynn):

Yep, you got it.